OpenStudy (anonymous):

@soccerman Desiree put $175 into a CD that pays 6% interest compounded semiannually. According to the rule of 72, approximately how long will it take for her money to double? A. 8 years B. 12 years C. 18 years D. 16 years

OpenStudy (anonymous):

@lasttccasey

OpenStudy (lasttccasey):

You must use A=P(1+r/n)^(nt), where A = total accrued amount, P is principle, r is the rate, n is the number of compounds annually, and t is time in years. Using this you find t = 11.724 Years so the closest answer is 12 Years.

OpenStudy (jack1):

@lasttccasey whats the rule of 72?

OpenStudy (lasttccasey):

That's basically telling you the rate needs to be divided by the compounded amount which is semiannually so its 2 in this case. Also that's the letter 'n' in the formula above. According to Wiki: http://en.wikipedia.org/wiki/Rule_of_72

OpenStudy (jack1):

think i get it better number of periods = 72/interest rate per period = 72/3 = 24 so 24 periods semiannually = 12 years?

OpenStudy (lasttccasey):

Actually the true formula is ln(2)/ln(1+r) = T in years. I'm not very familiar with this stuff but doing that formula you get 11.90 years basically. I'm studying mechanical engineering so this is a little outside my area of practice lol

OpenStudy (jack1):

ahha, cool, cheers for the help tho man, gonna read up on it now

OpenStudy (lasttccasey):

Nice, well good luck and you're welcome.